# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 2.2. A geometric explanation
We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{vi}}$ are the same as those in the domain $D_{11}$, and the remaining cases are similar.
If $c_{ijk}$ is in the domain $D_{11}$, then $q_{i}<0,h_{j}<0$ and the rest is positive by Table 1. Suppose $c_{ijk}$ is in the domain $D_{\mathrm{vi}}$. By Lemma 2.5, the vectors $c_{ijk},c_{ij},v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ij}\in E_{ij}$ is not in $\{ij\}$. Furthermore, since $c_{ijk}$ is in $D_{\mathrm{vi}}$, then $c_{ij}$ is on the opposite side of $j$ as $i$ and hence $d_{ij}<0,d_{ji}>0$. Similarly, we have $d_{ik}<0,d_{ki}>0,d_{jk}>0,d_{kj}>0$. By Proposition 2.14, we have $h_{ij}<0,h_{jk}>0,h_{ik}>0$. By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{jk},v_{i}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, then $c^{\prime}_{jk}\in E^{\prime}_{jk}$ is not in $i$. Furthermore, since $c_{ijk}$ is in $D_{\mathrm{vi}}$, then $c^{\prime}_{jk}$ is on the opposite side of $\{ik\}$ as $\{ij\}$ and hence $\theta_{jk}<0,\theta_{kj}>0$. Similarly, we have $\theta_{ji}<0,\theta_{ij}>0,\theta_{ik}>0,\theta_{ki}>0$. By Proposition 2.14, we have $q_{ij}>0,q_{jk}<0,q_{ik}>0$. Q.E.D.
By direct calculations, we can get the following results. As the proofs of the Lemma 2.10, Proposition 2.12 and Proposition 2.14 are too long, we defer them to Appendix A. For simplicity, we set $h_{t}=h_{rs},\ q_{t}=q_{rs}$ for $\{r,s,t\}=\{i,j,k\}$.
@@
As a direct corollary of Lemma 2.10, we have
@@
If $c_{ijk}$ is time-like, then the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
@@
If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ and $q$, and $\theta$ and $d$.
If $c_{ijk}$ lies in the domains $D_{\mathrm{I}},D_{\mathrm{III}},D_{\mathrm{IV}}$ or $D_{\mathrm{VI}}$, we have similar propositions. Using these propositions, we can prove Lemma 2.8 and further obtain Table 1. Moreover, from Table 1, we see that $c_{ijk}$ lies in the open domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}}$ or $D_{\mathrm{VI}}$ if and only if one of $q_{i},q_{j},q_{k},h_{i},h_{j},h_{k}$ is negative and the others are positive. Hence, these propositions can be written as a unified form.
@@
If $h_{r}<0$, then we swap $h,\ \theta$ and $q,\ d$.
@@
Similarly, if $c_{ijk}$ lies in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}}$ or $D_{\mathrm{v}}$, we have the corresponding propositions. Using these propositions, we can prove Lemma 2.9 and further obtain Table 1. Moreover, from Table 1, we see that $c_{ijk}$ lies in the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}}$ or $D_{\mathrm{vi}}$ if and only if $q_{r}<0,h_{s}<0$ and the rest are positive. Hence, these propositions can also be written as a unified form.
@@
Let $c_{ijk}\in\mathbb{R}^{3}$ lie in one of the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$. If $q_{r}<0,h_{s}<0$, then
@@
As a direct corollary of Lemma 2.13 and Lemma 2.15, we have
@@
If $c_{ijk}$ is space-like, then the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
@@
For any right-angled hyperbolic hexagon $\{ijk\}\in F$, the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
@@
If $c_{ijk}$ is time-like, the conclusion follows from Corollary 2.11. If $c_{ijk}$ is space-like, the conclusion follows from Corollary 2.16. If $c_{ijk}$ is light-like, the conclusion follows from continuity by moving the center to the light cone. Q.E.D.
@@ 2.3. Variational formulas of generalized angle
We divide Theorem 1.12 into Lemma 2.18, Lemma 2.20 and Lemma 2.22, and prove them respectively.
@@
For a right-angled hyperbolic hexagon $\{ijk\}\in F$, we have
@@
where $\beta=1$ if $c_{ijk}$ is time-like and $\beta=-1$ if $c_{ijk}$ is space-like. If $c_{ijk}$ is light-like, we interpret the formula as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
@@
Without loss of generality, we assume $||v_{i}||=||v_{j}||=||v_{k}||=1$ and $||c_{ijk}||^{2}=\pm 1$. For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r},\ h_{r}=h_{st},\ q_{r}=q_{st}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. By generalized hyperbolic cosine law, i.e., $\cosh l_{r}=\cosh\theta_{r}\sinh l_{s}\sinh l_{t}-\cosh l_{s}\cosh l_{t}$, we have
@@
where (13) is used in the forth line. Set
@@
To calculate $L_{1}$ and $L_{2}$, we first give some formulas. Please refer to Figure 3. Applying Proposition 2.1 to the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$ gives
(20) $v_{j}\ast c_{ijk}=-(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})=\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk}),$
where $v_{j}\ast c_{ij}=-\sinh d_{ji}$ follows from Proposition 2.2 (ii). Similarly, in the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have
@@
By Proposition 2.2 $\mathrm{(iii)}$, we have $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{jk}||(-\cosh\theta_{ik}).$ Moreover,
$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=(c_{jk}\ast c_{ijk})||v_{j}||^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{jk})=(c_{jk}\ast c_{ijk})\cdot\cosh^{2}d_{jk},$
where (21) is used in the second equality. Since $||v_{j}\otimes c_{jk}||^{2}=(v_{j}\ast c_{jk})^{2}+1=\cosh^{2}d_{jk}$, then
(23) $\cosh\theta_{ik}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})}{||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{jk}||}=-\frac{(c_{jk}\ast c_{ijk})\cdot\cosh d_{jk}}{||v_{j}\otimes c_{ijk}||}.$
Similarly, $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{ij}||(-\cosh\theta_{ki}).$ Moreover,
$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=(c_{ij}\ast c_{ijk})||v_{j}||^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{ij})=(c_{ij}\ast c_{ijk})\cdot\cosh^{2}d_{ji},$
where (20) is used in the second equality. Since $||v_{j}\otimes c_{ij}||^{2}=(v_{j}\ast c_{ij})^{2}+1=\cosh^{2}d_{ji}$, then
(24) $\cosh\theta_{ki}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})}{||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{ij}||}=-\frac{(c_{ij}\ast c_{ijk})\cdot\cosh d_{ji}}{||v_{j}\otimes c_{ijk}||}.$
@@
(25) $||v_{j}\otimes c_{ijk}||^{2}=(v_{j}\otimes c_{ijk})*(v_{j}\otimes c_{ijk})=(v_{j}\ast c_{ijk})^{2}-||c_{ijk}||^{2},$
@@
(26) $||c_{ij}\otimes c_{ijk}||^{2}=(c_{ij}\otimes c_{ijk})*(c_{ij}\otimes c_{ijk})=(c_{ij}\ast c_{ijk})^{2}+||c_{ijk}||^{2}.$
@@
$=$ $\frac{1}{||v_{j}\otimes c_{ijk}||^{2}}[(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}-||v_{j}\otimes c_{ijk}||^{2}]$
$=$ $\frac{||c_{ij}\otimes c_{ijk}||^{2}}{||v_{j}\otimes c_{ijk}||^{2}},$
@@
Now we begin to calculate $L_{1}$ and $L_{2}$ in (18).
@@
$=$ $-\sinh\theta_{ki}\cdot\sinh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}\cdot\frac{||v_{j}\otimes c_{ijk}||}{c_{ij}\ast c_{ijk}}\cdot\frac{c_{ij}\ast c_{ijk}}{||v_{j}\otimes c_{ijk}||}$
$=$ $\sinh\theta_{ki}\sinh\theta_{ik}\cdot\frac{||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\cosh\theta_{ki},$
where (24) is used in the last line.
@@
$=$ $\frac{\cosh d_{jk}\cdot(c_{jk}\ast c_{ijk})}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})}+\frac{(c_{ij}\ast c_{ijk})\cosh d_{ji}}{||v_{j}\otimes c_{ijk}||}\cdot\cosh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}$
$=$ $\frac{\cosh\theta_{ik}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot||v_{j}\otimes c_{ijk}||}\big{(}-||v_{j}\otimes c_{ijk}||^{2}+(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}\big{)}$
$=$ $\frac{\cosh\theta_{ik}\cdot||c_{ij}\otimes c_{ijk}||^{2}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot||v_{j}\otimes c_{ijk}||}$
$=$ $\frac{\cosh\theta_{ik}\sinh^{2}\theta_{ki}||v_{j}\otimes c_{ijk}||}{\sinh d_{ji}(c_{ij}\ast c_{ijk})},$
where (22) and (24) are used in the second line, (23) is used in the third line, (27) is used in the forth line, and (28) is used in the last line. Therefore, the formula (19) can be written as
@@
$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot(\cosh\theta_{ik}\sinh\theta_{ki}+\sinh\theta_{ik}\cosh\theta_{ki})$
$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\sinh\theta_{j}$
$=$ $\frac{1}{\sinh l_{k}\sinh d_{ji}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{c_{ij}\ast c_{ijk}}.$
(i): Suppose $c_{ijk}$ is time-like, then $||c_{ijk}||^{2}=-1$ and $v_{j}*c_{ijk}=-\sinh q_{j}$. By (25), we have $||v_{j}\otimes c_{ijk}||=\cosh q_{j}$. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\cosh q_{j}=\sinh h_{k}$ by Lemma 2.10. Since $c_{ij}\ast c_{ijk}=-\cosh h_{k}$, then

### v3

@@ 2.2. A geometric interpretation
We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{vi}}$ coincide with those in the domain $D_{11}$, and the remaining cases follow analogously.
If $c_{ijk}\in D_{11}$, then $q_{jk}<0$, $h_{ik}<0$, and the rest are positive by Table 1. Suppose $c_{ijk}\in D_{\mathrm{vi}}$. By Lemma 2.5, the vectors $c_{ijk}$, $c_{ij}$, and $v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Since $c_{ijk}\notin v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it follows that $c_{ij}\notin\{ij\}$. Furthermore, since $c_{ij}$ lies on the opposite side of $j$ as $i$, it follows that $d_{ij}<0$ and $d_{ji}>0$. Similarly, $d_{ik}<0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$. By Proposition 2.13, we have $h_{ij}<0$, $h_{jk}>0$, and $h_{ik}>0$. By Lemma 2.6, the vectors $c_{ijk}$, $c^{\prime}_{jk}$, and $v_{i}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Since $c_{ijk}\notin v_{i}v_{j}v_{k}$, it follows that $c^{\prime}_{jk}\notin i$. Furthermore, since $c^{\prime}_{jk}$ lies on the opposite side of $\{ik\}$ as $\{ij\}$, it follows that $\theta_{jk}<0$ and $\theta_{kj}>0$. Similarly, $\theta_{ji}<0$, $\theta_{ij}>0$, $\theta_{ik}>0$, and $\theta_{ki}>0$. By Proposition 2.13, we have $q_{ij}>0$, $q_{jk}<0$, and $q_{ik}>0$. Q.E.D.
By direct calculations, we obtain the following results. As the proofs of Lemma 2.9, Proposition 2.11 and Proposition 2.13 are too long, we defer them to Appendix A. For simplicity, we set $h_{t}=h_{rs}$ and $q_{t}=q_{rs}$ for $\{r,s,t\}=\{i,j,k\}$.
@@
As a direct corollary of Lemma 2.9, we have the following result.
@@
If $c_{ijk}$ is time-like, then the signs of $q_{r}$, $d_{rt}$, and $d_{rs}$ coincide, and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ coincide.
@@
If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ with $q$, and $\theta$ with $d$.
If $c_{ijk}$ lies in the domains $D_{\mathrm{I}},D_{\mathrm{III}},D_{\mathrm{IV}}$ or $D_{\mathrm{VI}}$, analogous propositions hold. These propositions enable us to prove Lemma 2.7 and further derive Table 1. Moreover, Table 1 shows that $c_{ijk}$ lies in the open domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}}$ or $D_{\mathrm{VI}}$ if and only if one of $q_{i},q_{j},q_{k},h_{i},h_{j},h_{k}$ is negative and the rest are positive. Consequently, these propositions can be written in a unified form.
@@
If $h_{r}<0$, then we swap $h$ with $q$, and $\theta$ with $d$.
@@
Similarly, if $c_{ijk}$ lies in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}}$ or $D_{\mathrm{v}}$, corresponding propositions hold. These propositions allow us to prove Lemma 2.8 and further derive Table 1. Moreover, Table 1 shows that $c_{ijk}$ lies in the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}}$ or $D_{\mathrm{vi}}$ if and only if $q_{r}<0$, $h_{s}<0$, and the rest are positive. Consequently, these propositions also admit a unified form.
@@
Let $c_{ijk}\in\mathbb{R}^{3}$ lie in one of the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$. If $q_{r}<0$ and $h_{s}<0$, then
@@
As a direct corollary of Lemma 2.12 and Lemma 2.14, we have the following result.
@@
If $c_{ijk}$ is space-like, then the signs of $q_{r}$, $d_{rt}$, and $d_{rs}$ coincide, and the signs of $h_{r}$, $\theta_{rs}$, and $\theta_{rt}$ coincide.
@@
For any right-angled hyperbolic hexagon $\{ijk\}\in F$, the signs of $q_{r}$, $d_{rt}$, and $d_{rs}$ coincide, and the signs of $h_{r}$, $\theta_{rs}$, and $\theta_{rt}$ coincide.
@@
If $c_{ijk}$ is time-like, the conclusion follows from Corollary 2.10. If $c_{ijk}$ is space-like, the conclusion is derived from Corollary 2.15. If $c_{ijk}$ is light-like, the conclusion holds by continuity upon moving $c_{ijk}$ to the light cone. Q.E.D.
@@ 2.3. Variational formulas of generalized angle
We decompose Theorem 1.11 into Lemma 2.17, Lemma 2.19, and Lemma 2.21, and prove them respectively.
@@
For any right-angled hyperbolic hexagon $\{ijk\}\in F$, we have
@@
where $\beta=1$ if $c_{ijk}$ is time-like, and $\beta=-1$ if $c_{ijk}$ is space-like. If $c_{ijk}$ is light-like, the formula is interpreted as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
@@
Without loss of generality, we assume that $\|v_{i}\|=\|v_{j}\|=\|v_{k}\|=1$ and $\|c_{ijk}\|^{2}=\pm 1$. For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r},\ h_{r}=h_{st},\ q_{r}=q_{st}$, and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. By generalized hyperbolic cosine law, i.e., $\cosh l_{r}=\cosh\theta_{r}\sinh l_{s}\sinh l_{t}-\cosh l_{s}\cosh l_{t}$, we derive
@@
where (13) is used in the fourth line. Define
@@
To compute $L_{1}$ and $L_{2}$, some preliminary results are required. We begin by referring to Figure 3. By Proposition 2.2 (ii), we have $v_{j}\ast c_{ij}=+\sinh d_{\mathbb{H}}(v_{j}^{\bot},c_{ij})=+\sinh(-d_{ji})=-\sinh d_{ji}$, or $v_{j}\ast c_{ij}=-\sinh d_{\mathbb{H}}(v_{j}^{\bot},c_{ij})=-\sinh(+d_{ji})=-\sinh d_{ji}$. In both cases, it follows that $v_{j}\ast c_{ij}=-\sinh d_{ji}$. Applying Proposition 2.1 to the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$ yields
(19) $v_{j}\ast c_{ijk}=-(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})=\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk}).$
Similarly, in the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we derive
@@
By Proposition 2.2 (iii), we have $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=\|v_{j}\otimes c_{ijk}\|\cdot\|v_{j}\otimes c_{jk}\|(-\cosh\theta_{ik}).$ Moreover, direct calculation yields
$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=(c_{jk}\ast c_{ijk})\|v_{j}\|^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{jk})=(c_{jk}\ast c_{ijk})\cdot\cosh^{2}d_{jk},$
where (20) is used in the second equality. Since $\|v_{j}\otimes c_{jk}\|^{2}=(v_{j}\ast c_{jk})^{2}+1=\cosh^{2}d_{jk}$, it follows that
(22) $\cosh\theta_{ik}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})}{\|v_{j}\otimes c_{ijk}\|\cdot\|v_{j}\otimes c_{jk}\|}=-\frac{(c_{jk}\ast c_{ijk})\cdot\cosh d_{jk}}{\|v_{j}\otimes c_{ijk}\|}.$
Similarly, $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=\|v_{j}\otimes c_{ijk}\|\cdot\|v_{j}\otimes c_{ij}\|(-\cosh\theta_{ki}).$ Moreover,
$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=(c_{ij}\ast c_{ijk})\|v_{j}\|^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{ij})=(c_{ij}\ast c_{ijk})\cdot\cosh^{2}d_{ji},$
where (19) is used in the second equality. Since $\|v_{j}\otimes c_{ij}\|^{2}=(v_{j}\ast c_{ij})^{2}+1=\cosh^{2}d_{ji}$, it follows that
(23) $\cosh\theta_{ki}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})}{\|v_{j}\otimes c_{ijk}\|\cdot\|v_{j}\otimes c_{ij}\|}=-\frac{(c_{ij}\ast c_{ijk})\cdot\cosh d_{ji}}{\|v_{j}\otimes c_{ijk}\|}.$
@@
(24) $\|v_{j}\otimes c_{ijk}\|^{2}=(v_{j}\otimes c_{ijk})*(v_{j}\otimes c_{ijk})=(v_{j}\ast c_{ijk})^{2}-\|c_{ijk}\|^{2},$
@@
(25) $\|c_{ij}\otimes c_{ijk}\|^{2}=(c_{ij}\otimes c_{ijk})*(c_{ij}\otimes c_{ijk})=(c_{ij}\ast c_{ijk})^{2}+\|c_{ijk}\|^{2}.$
@@
$=\frac{1}{\|v_{j}\otimes c_{ijk}\|^{2}}[(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}-\|v_{j}\otimes c_{ijk}\|^{2}]$
$=\frac{\|c_{ij}\otimes c_{ijk}\|^{2}}{\|v_{j}\otimes c_{ijk}\|^{2}},$
@@
We are now ready to compute $L_{1}$ and $L_{2}$ in (17). For $L_{2}$, we proceed as follows:
@@
$=-\sinh\theta_{ki}\cdot\sinh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}\cdot\frac{\|v_{j}\otimes c_{ijk}\|}{c_{ij}\ast c_{ijk}}\cdot\frac{c_{ij}\ast c_{ijk}}{\|v_{j}\otimes c_{ijk}\|}$
$=\sinh\theta_{ki}\sinh\theta_{ik}\cdot\frac{\|v_{j}\otimes c_{ijk}\|}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\cosh\theta_{ki},$
where (23) is used in the last line. For $L_{1}$, we have
@@
$=\frac{\cosh d_{jk}\cdot(c_{jk}\ast c_{ijk})}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})}+\frac{(c_{ij}\ast c_{ijk})\cosh d_{ji}}{\|v_{j}\otimes c_{ijk}\|}\cdot\cosh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}$
$=\frac{\cosh\theta_{ik}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot\|v_{j}\otimes c_{ijk}\|}\big{(}-\|v_{j}\otimes c_{ijk}\|^{2}+(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}\big{)}$
$=\frac{\cosh\theta_{ik}\cdot\|c_{ij}\otimes c_{ijk}\|^{2}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot\|v_{j}\otimes c_{ijk}\|}$
$=\frac{\cosh\theta_{ik}\sinh^{2}\theta_{ki}\|v_{j}\otimes c_{ijk}\|}{\sinh d_{ji}(c_{ij}\ast c_{ijk})},$
where (21) and (23) are used in the second line, (22) is used in the third line, (26) is used in the fourth line, and (27) is used in the last line. Substituting $L_{1}$ and $L_{2}$ into (18) yields
@@
$=\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot(\cosh\theta_{ik}\sinh\theta_{ki}+\sinh\theta_{ik}\cosh\theta_{ki})$
$=\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\sinh\theta_{j}$
$=\frac{1}{\sinh l_{k}\sinh d_{ji}}\cdot\frac{\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|}{c_{ij}\ast c_{ijk}}.$
(i): Assume that $c_{ijk}$ is time-like. Then $\|c_{ijk}\|^{2}=-1$ and $v_{j}*c_{ijk}=-\sinh q_{j}$. By (24), we have $\|v_{j}\otimes c_{ijk}\|=\cosh q_{j}$. Furthermore, by Lemma 2.9, we obtain $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\cosh q_{j}=\sinh h_{k}$. Since $c_{ij}\ast c_{ijk}=-\cosh h_{k}$, it follows that